WileyPLUS Problem 8-5 a-b (Part Level Submission) Company A had purchased machinery 3 years ago at the beginning of 2013 andPart (b) only
Answer #1

Depreciation under Straight line method = (Cost - Residual value) / Estimated useful life

= ($46,000 - $3,800) / 5

= $8,440

Accumulated depreciation for 3 years = $8,440 * 3 = $25,320

Book value = $46,000 - $25,320 = $20,680

Remaining useful life = 7 - 3 = 4 years

Revised annual depreciation = (Book value - Residual value) / Remaining useful life

= ($20,680 - $1,200) / 4

= $4,870